A forester mixes gasoline and oil to make 2 gallons of mixture for his two-cycle chain-saw engine. This mixture is 32 parts gasoline and 1 part two-cycle oil. How much gasoline must be added to bring the mixture to 40 part gasoline and 1 part oil?
I tried to make a chart to list everything I knew from the problem and see how I could solve it, but I'm not sure where to start from.

Answer by ptaylor(2048)

Let x=amount of gas that needs to be added
Now we know that the amount of oil in the 32-1 mixture is the same as the amount of oil in the 40-1 mixture (since we neither add or take away any oil); so:
2*(1/32)=1/40(2+x) get rid of parens
1/16=1/20 + x/40 (LCD=80) multiply each term by 80
5=4+2x subtract 4 from each side
5-4=2x divide by 2
x=1/2 gal amount of gas that needs to be added
CK
(1/32)*2=(1/40)*2.5
1/16=1/16
Hope this helps------ptaylor